OpenStudy (anonymous):

Identify the horizontal asymptote of f(x) = 6x -7 / 11x+8

OpenStudy (jhannybean):

To find the horizontal asymptote compare the variable with the highest degree in the numerator to the variable with the highest degree in the denominator.

OpenStudy (jhannybean):

\[\large f(x) = \frac{ 6\color{red}x-7}{11\color{red}x+8}\] The variables are both to the power of one, therefore your aymptote is the result of their coefficients.

OpenStudy (anonymous):

Which would end up being HA=6/11

OpenStudy (jhannybean):

\[\large f(x) = \frac{ \color{blue}6\color{red}{x^1}-7}{\color{blue}11\color{red}{x^1}+8}\]

OpenStudy (anonymous):

thankyou guys ! :)

OpenStudy (jhannybean):

No problemo.

OpenStudy (anonymous):

and thankyou for you explanation jhannybean